OpenStudy (anonymous):

(10x-3)(4x+5) A.14x^2+62x-15 B.40x^2+38x-15 C.40x^2+38x+2 D.40x^2+62x-15

OpenStudy (skullpatrol):

have you heard of FOIL?

OpenStudy (anonymous):

FOIL. First Out Inner Last

OpenStudy (anonymous):

? dont know how to do that

OpenStudy (skullpatrol):

have you heard of the distributive property?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

help

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (skullpatrol):

Every time you multiply two numbers you are using this^ property

OpenStudy (skullpatrol):

$$\Huge (10x-3)(4x+5) $$ $$\Huge=(10x-3)(4x)+(10x-3)(5) $$ $$\Huge=40x^2-12x+50x-15$$ $$\Huge =40x^2+38x-15$$

OpenStudy (anonymous):

Thank You

OpenStudy (skullpatrol):

Thanks for asking :)